# Protocol as Poetry: Case Study on Pak’s Protocol Arts

### v1

@@ A.7. Merge
The launch of Merge saw an enormous turnout, and this critical mass of participants created a unique collector community bonded by the experiment. Immediately, Merge holders began comparing the sizes of their masses, fostering playful rivalries between “whales” (those with very large masses) and everyday collectors. There was a leaderboard mentality; some individuals who acquired huge quantities of mass became minor celebrities in the community, their holdings seen as a bold statement of support for Pak’s vision. On secondary markets, novel behaviors emerged: because any purchase that combined two masses was irreversible, some collectors were hesitant to buy an additional Merge token unless it was larger than the one they already possessed, to avoid ‘losing’ their smaller piece in a merge. This led to an unusual trading strategy where sometimes whole wallets (with a Merge NFT inside) were sold peer-to-peer, just to preserve the token’s separate identity. The theoretical endgame of one entity eventually merging all masses into “The One” became a topic of both amusement and intrigue, serving as a metaphor for consolidation of power or unity of art. Meanwhile, the sheer fact that Merge had broken the world record for an art sale – not by one person’s expenditure but by the collective effort of nearly 29,000 buyers – instilled a sense of communal pride. This remains a profound aspect of the piece’s narrative: it proved that a decentralized network of collectors could, together, make art history. Over time, as some merges did occur and the token count slowly diminished, the community continued to monitor and discuss these developments, treating the artwork as an ongoing story they were all part of. Merge thus nurtured both camaraderie and competition among its participants, encapsulating a microcosm of the crypto community’s spirit.
@@
On a conceptual level, Merge operates as a poetic exploration of unity and the power of the many. It turned the act of collecting into a narrative of convergence: each individual NFT was not just an endpoint but a piece of a larger potential composition. There is an inherent poetic image in thousands of separate collectors holding what could be fragments of one singular artwork – it’s as if a digital star was shattered and distributed, with the ever-present possibility of reassembly. This resonates with ideas of community in the crypto space: Merge implicitly asks whether the true value of an artwork might reside in shared ownership and collective action rather than exclusivity. The artwork’s very existence and record-breaking price were a testament to collaboration (a crowd of buyers rather than a single bidder), making it a celebration of a decentralized patronage model. Additionally, the merging mechanic itself can be seen as a commentary on accumulation and synthesis: smaller masses being absorbed by bigger ones evokes phenomena in economics and nature alike, delivered here as a voluntary game. By provoking participants to consider merging or resisting merges, Pak surfaced questions about competition versus cooperation and about whether art is better experienced as a multitude or as a unity. Ultimately, Merge’s broader impact lies in its demonstration that scarcity and unity can be artistically intertwined: it delivered a visual and participatory poem about how the many can become one. In doing so, it expanded the discourse of digital art to include not just the creation of images, but the creation of new social contracts and economic structures as vehicles for meaning.
@@
Merge was sold on 2–4 December 2021 and drew 28,983 buyers who purchased 312,686 total mass units . The sale grossed \$91.8 million , making it (at that time) the highest total for any NFT artwork and thrusting Pak above Jeff Koons as the priciest living artist by primary market sales. The average collector bought about 10.8 mass units. Initially, 28,983 Merge tokens existed (one per buyer). Through secondary-market activity and merging, the supply has been decreasing: one year post-sale, the number of distinct Merge tokens had dropped to around 27,000 as consolidations happened. The largest token (Alpha) amassed 933,878.2 mass (some collectors found ways to add mass beyond the sale via special mechanics or bonuses), and the next few largest were orders of magnitude smaller, highlighting a steep consolidation curve. The smart contract ensured that a Merge token’s mass count and visual size update in real-time with each merge; Nifty Gateway’s interface had to be adapted to handle these dynamic NFTs. By January 2022, Merge tokens had done over \$100 million in secondary trading volume as collectors continued to jockey for position. Importantly, no single entity has (yet) merged all tokens – tens of thousands of decentralized pieces remain, meaning Merge lives on as a plural artwork owned by many, with the theoretical possibility that it could one day coalesce further (Thomas, 2021).
@@ A.8. Censored
Figure 11. ”Censored is a Collection By Pak and Assange and You”. Censored (2022); The top figure shows the soul-bound censored NFT, which remains untransferable until Assange’s release. The bottom figure shows the uncensored version, which becomes transferable / tradable after Assange’s release.
@@
Censored (February 2022) is a two-part collaborative artwork by Pak and WikiLeaks founder Julian Assange that confronts issues of censorship, free speech, and the power of the collective. Conceived while Assange was imprisoned, the project’s central question is how information and truth can be either suppressed or liberated in the digital age. The first part, Clock, is a one-of-a-kind dynamic NFT that simply displays a count of the number of days Assange has been in custody – an ever-increasing tally that starkly visualizes the passage of time lost to censorship and incarceration (Cen, 2022). The second part was a dynamic open edition often referred to as Censored (later “Uncensored”), which invited anyone to anonymously submit a text message. Each submission was minted as an NFT where the message initially appeared as blacked-out blocks (as if redacted), but later these NFTs were revealed to show the full message – creating a permanent, public ledger of uncensored speech on the blockchain . Together, these components embody a powerful concept: Censored is both a protest and a platform, highlighting the plight of Assange specifically and the principle of uncensored expression universally. By design, the project blurred the line between artist and participant, making “you” – the public – a co-creator of the open edition piece, echoing the idea that free expression is a collective, democratic effort.
@@
The technical structure of Censored mirrored its conceptual duality. Clock was implemented as a dynamic NFT: its content updates once per day to increment the count of Assange’s imprisonment. This likely involves either an on-chain calculation of days since a start date or an off-chain oracle pushing daily updates – either way, the piece is programmatically tied to the real-world passage of time and will continue to tick until Assange’s status changes. The open edition was dynamic in a different sense: when each Censored NFT was first minted, the text submitted by the user was stored (on-chain or via metadata) but visually represented in a redacted style (for example, as black bars or scrambled characters). Then, at a predetermined time, a contract function executed to “uncensor” all these tokens, perhaps by updating a base image or revealing a hidden layer in the metadata so that the true text became visible. Each open-edition NFT is thus unique in message but identical in having undergone the same transformation from obscured to revealed. The fundraising mechanism was straightforward: the sale smart contract directed the incoming Ether to designated wallets associated with Assange’s defense fund and other aligned charities (indeed, a portion of the 671 ETH raised from the open edition was later donated to support journalistic freedom and even humanitarian aid ). The dynamics here combined artistic intent with real-world impact: the more people voiced themselves (minted NFTs) and the higher the DAO bid for Clock, the more tangible support was generated for the cause. In a sense, the blockchain functions as both canvas and conduit in Censored, ensuring that the art’s message (the imperative of free expression) is inseparable from the action it precipitates (material aid for a free press).
@@
The participatory element of Censored was striking. For Clock, participation took the form of collective fundraising: an Assange-supporter DAO (decentralized autonomous organization) pooled resources from over 10,000 people to bid on and ultimately win the piece for 16,593 ETH (about \$52.8 million) , making each contributor a part-owner of the artwork and massively amplifying its political message. For the open edition, Pak opened a web portal on February 7, 2022, where anyone could type a short message and mint it as an NFT. There were minimal barriers – aside from blockchain transaction fees – which allowed thousands of ordinary people worldwide to inscribe their uncensored thoughts onto the blockchain. During the mint period, participants wrote messages ranging from political statements (“FREE ASSANGE NOW”) to personal reflections, fully aware that they would initially be recorded in a censored (illegible) state. After the mint closed, Pak triggered a global reveal: all the NFTs’ images were updated to show the actual text of the messages, symbolizing voices breaking through suppression. In this way, the audience’s role wasn’t just to observe but to speak and be memorialized as part of the art. Additionally, by buying these open-edition NFTs (each priced modestly), participants contributed to a fundraiser – all proceeds were directed to Assange’s legal defense and pro-freedom organizations – blending activism with artistic engagement . In essence, Censored turned its audience into an active assembly of publishers and protesters, demonstrating the principle that in a decentralized network, everyone can raise their voice and collectively bolster a cause.
@@
In Censored, Pak acted not only as an artist but also as a facilitator of collective expression and activism. He and Assange’s team set the stage – defining the parameters of the open edition (time frame, interface for input, price) and crafting the Clock NFT – but then handed the mike to the public. For the open edition, Pak exercised restraint in content moderation: aside from perhaps filtering out nothing (to align with the absolutist free speech stance), he allowed minters to submit any text, which was a radical trust in the community. However, Pak retained control over the overall execution: he decided when the reveal would happen and ensured it occurred uniformly for everyone, and he controlled the smart contract that forwarded funds to the intended beneficiaries. With Clock, once it was launched, the control shifted to the AssangeDAO (as the new owners) in terms of display and eventual fate of the piece, although the daily increment feature was baked into the token itself. In essence, Pak’s control was about creating a secure and open channel for expression and then stepping back at the right moments. The success of Censored depended on this light-touch approach; too much control (e.g., censoring the messages or manipulating the outcomes) would have undermined the trust and authenticity the piece needed to resonate. By structuring it as he did, Pak underscored his role as an enabler — he controlled the framework but not the voices within, embodying the artwork’s ethos in the very way he managed it.
Collective Emerging Behavior
@@
The Censored project galvanized a uniquely blended community of art enthusiasts, cypherpunks, transparency advocates, and supporters of Assange. The formation of AssangeDAO itself was a landmark event: thousands of strangers coordinated in a matter of days on chat platforms, raised tens of millions in cryptocurrency, and collectively won the Clock auction, an unprecedented feat of decentralized crowd-funding . This feat gave participants a profound sense of accomplishment – each member of the DAO owns a fractional piece of Clock via tokens, making Clock a collectively owned symbol of protest. In parallel, the open edition saw likely tens of thousands of NFTs minted (the exact number of messages minted was significant, suggesting a broad participation from the crypto community and beyond). Upon the reveal, a new collective experience began: people scoured the now-uncensored messages, sharing poignant or powerful quotes they found among the NFTs. The collection became a digital chorus of statements about freedom, justice, and personal sentiments, and owning one meant being part of that chorus. A notable emergent behavior was the way traditional lines blurred: activists became art collectors by minting NFTs, and NFT collectors became activists by joining the political messaging. The Censored Discord and other forums buzzed with discussion not just about the art’s value, but about the cause it supported and the stories behind individual messages. This convergence of normally separate communities (art, crypto, activism) was itself a product of the project. In the aftermath, the collective energy didn’t dissipate immediately: the DAO refocused on continued advocacy for Assange, and many open-edition holders continued to promote the messages they had minted, effectively using their NFTs as badges of alignment with the free speech movement. Censored thus forged an alliance of convenience into a lasting community centered on principle and expression.
@@
Censored stands out as a poignant union of art and activism, transforming the cold mechanics of blockchain into an emotional and political narrative. The poetic substance of the work arises from its contrasts: a single, ever-ticking clock counting the days of one man’s silencing, and a sea of uncensorable messages giving voice to thousands. The Clock NFT is minimalistic yet profound – each passing day it displays is a testimony to injustice, a visual poem of waiting and resilience. The open edition, once uncensored, became an anthology of global voices; its poetry is both literal (many entries were written with genuine passion or wit) and metaphorical, in that it turned censorship on its head – what was once hidden is now immortalized in public view. There is a strong sentiment that Censored wasn’t just documenting a moment in crypto art, but a moment in history: it bridged the gap between the decentralized art world and real-world social issues, demonstrating that the blockchain community could rally around something deeply human. The broader impact of Censored lies in its demonstration of how digital art can be immediately socially relevant. It set a precedent for using NFTs as a medium of protest and fundraising, expanding the notion of what kind of statements can be made – and preserved – through art. In summary, Censored made poetic meaning through direct action: it is art as a verb, an event in which aesthetic expression, technological infrastructure, and moral conviction converged to powerful effect.
@@
The Clock NFT sold on 9 February 2022 for 16,593 ETH (approximately \$52.8 million) , bought by AssangeDAO, a collective of 10,000+ members pooling funds . This made it one of the most expensive NFTs ever and provided a huge donation to Assange’s legal defense. The open edition X/X minted 29,766 NFTs over its 48-hour run , raising 671 ETH in voluntary contributions (about \$2.1 million) for pro-freedom organizations chosen by Pak and Assange . Each of these message NFTs remains non-transferable and will only become tradeable upon Assange’s release . Over 17,000 unique Ethereum addresses participated in the mint (many contributed multiple messages via additional wallets). The largest donation from a single minter was 50 ETH (showing some used the mint as a way to donate significantly). Post-drop, AssangeDAO has kept Clock on display via a fractionalized ownership governance token (representing membership shares in the DAO). censored stands as a record-breaking instance of political fundraising through digital art, and a case where an NFT project directly engaged international press and communities far beyond the traditional art world.

### v2

@@ A.7. Merge
The launch of Merge saw an enormous turnout, and this critical mass of participants created a unique collector community bonded by the experiment. Immediately, Merge holders began comparing the sizes of their masses, fostering playful rivalries between “whales” (those with very large masses) and everyday collectors. There was a leaderboard mentality; some individuals who acquired huge quantities of mass became minor celebrities in the community, their holdings seen as a bold statement of support for Pak’s vision. On secondary markets, novel behaviors emerged: because any purchase that combined two masses was irreversible, some collectors were hesitant to buy an additional Merge token unless it was larger than the one they already possessed, to avoid ‘losing’ their smaller piece in a merge. This led to an unusual trading strategy where sometimes whole wallets (with a Merge NFT inside) were sold peer-to-peer, just to preserve the token’s separate identity. The theoretical endgame of one entity eventually merging all masses into “The One” became a topic of both amusement and intrigue, serving as a metaphor for consolidation of power or unity of art. Meanwhile, the sheer fact that Merge had broken the world record for an art sale—not by one person’s expenditure but by the collective effort of nearly 29,000 buyers—instilled a sense of communal pride. This remains a profound aspect of the piece’s narrative: it proved that a decentralized network of collectors could, together, make art history. Over time, as some merges did occur and the token count slowly diminished, the community continued to monitor and discuss these developments, treating the artwork as an ongoing story they were all part of. Merge thus nurtured both camaraderie and competition among its participants, encapsulating a microcosm of the crypto community’s spirit.
@@
On a conceptual level, Merge operates as a poetic exploration of unity and the power of the many. It turned the act of collecting into a narrative of convergence: each individual NFT was not just an endpoint but a piece of a larger potential composition. There is an inherent poetic image in thousands of separate collectors holding what could be fragments of one singular artwork—it’s as if a digital star was shattered and distributed, with the ever-present possibility of reassembly. This resonates with ideas of community in the crypto space: Merge implicitly asks whether the true value of an artwork might reside in shared ownership and collective action rather than exclusivity. The artwork’s very existence and record-breaking price were a testament to collaboration (a crowd of buyers rather than a single bidder), making it a celebration of a decentralized patronage model. Additionally, the merging mechanic itself can be seen as a commentary on accumulation and synthesis: smaller masses being absorbed by bigger ones evokes phenomena in economics and nature alike, delivered here as a voluntary game. By provoking participants to consider merging or resisting merges, Pak surfaced questions about competition versus cooperation and about whether art is better experienced as a multitude or as a unity. Ultimately, Merge’s broader impact lies in its demonstration that scarcity and unity can be artistically intertwined: it delivered a visual and participatory poem about how the many can become one. In doing so, it expanded the discourse of digital art to include not just the creation of images, but the creation of new social contracts and economic structures as vehicles for meaning.
@@
Merge was sold on 2–4 December 2021 and drew 28,983 buyers who purchased 312,686 total mass units. The sale grossed \$91.8 million, making it (at that time) the highest total for any NFT artwork and thrusting Pak above Jeff Koons as the priciest living artist by primary market sales. The average collector bought about 10.8 mass units. Initially, 28,983 Merge tokens existed (one per buyer). Through secondary-market activity and merging, the supply has been decreasing: one year post-sale, the number of distinct Merge tokens had dropped to around 27,000 as consolidations happened. The largest token (Alpha) amassed 933,878.2 mass (some collectors found ways to add mass beyond the sale via special mechanics or bonuses), and the next few largest were orders of magnitude smaller, highlighting a steep consolidation curve. The smart contract ensured that a Merge token’s mass count and visual size update in real-time with each merge; Nifty Gateway’s interface had to be adapted to handle these dynamic NFTs. By January 2022, Merge tokens had done over \$100 million in secondary trading volume as collectors continued to jockey for position. Importantly, no single entity has (yet) merged all tokens—tens of thousands of decentralized pieces remain, meaning Merge lives on as a plural artwork owned by many, with the theoretical possibility that it could one day coalesce further (Thomas, 2021).
Annotation
Merge is the most ambitious and conceptually far-reaching work in the portfolio, serving as the definitive case for C1 (system-centric composition) and C2 (autonomous governance): once deployed, its self-merging deflationary logic operates without any artist intervention, and the artwork’s visual form (the size and color of each mass token) is entirely determined by on-chain state. After the initial open-edition sale, everything becomes an automatically self-governed universe in which the protocol alone dictates outcomes. C3 (distributed agency) reaches its largest scale here—28,983 buyers collectively sculpt the work’s final distribution—while C5 (economy-driven engagement) manifests not only in the frenzied mass-hoarding competition but in the emergence of the Internship DAO, a collectively organized effort to acquire the Alpha token that arose entirely from participant initiative rather than artist design. This emergent behavior is paradigmatic: user activity creates its own attentional dynamics, as though the work generates a distinct game that no one scripted. Every person, at every moment, faces the choice to sell or to acquire others’ tokens, and this perpetual decision space—playable at any time, but always with genuine economic incentive—transforms Merge into something closer to a persistent world than a static artwork. Merge is notable for the irreversibility of its dynamics: tokens can only grow through merging, never split, creating a one-directional “extinction game” that distinguishes it from the reversible or cyclical mechanics of burn.art and Lost Poets. This irreversibility gives Merge a distinctive temporal character (C4): the work can only become more concentrated over time, never return to its original state, ensuring that its aesthetic and economic trajectory is entropic and unrepeatable.
@@ A.8. Censored
Refer to caption A light gray background with the text “PAK WAS HERE” partially visible behind a thick black horizontal redaction bar that crosses through all three words, obscuring but not fully hiding the message.
Refer to caption A light gray background with the text “PAK WAS HERE” displayed clearly in bold black sans-serif capital letters, with no redaction bar, showing the fully revealed uncensored state of the NFT.
Figure 12. “Censored is a Collection by Pak and Assange and You.” Censored (2022). Top: a soul-bound message NFT in its censored state, with text obscured by a redaction bar—untransferable while Assange remains imprisoned. Bottom: the same NFT in its uncensored state after Assange’s release, revealing the full message and becoming freely transferable.
@@
Censored (February 2022) is a two-part collaborative artwork by Pak and WikiLeaks founder Julian Assange that confronts issues of censorship, free speech, and the power of the collective. Conceived while Assange was imprisoned, the project’s central question is how information and truth can be either suppressed or liberated in the digital age. The first part, Clock, is a one-of-a-kind dynamic NFT that simply displays a count of the number of days Assange has been in custody—an ever-increasing tally that starkly visualizes the passage of time lost to censorship and incarceration (Courage Foundation, 2022). The second part was a dynamic open edition often referred to as Censored (later “Uncensored”), which invited anyone to anonymously submit a text message. Each submission was minted as an NFT where the message initially appeared as blacked-out blocks (as if redacted), but later these NFTs were revealed to show the full message—creating a permanent, public ledger of uncensored speech on the blockchain. Together, these components embody a powerful concept: Censored is both a protest and a platform, highlighting the plight of Assange specifically and the principle of uncensored expression universally. By design, the project blurred the line between artist and participant, making “you”—the public—a co-creator of the open edition piece, echoing the idea that free expression is a collective, democratic effort.
@@
The technical structure of Censored mirrored its conceptual duality. Clock was implemented as a dynamic NFT: its content updates once per day to increment the count of Assange’s imprisonment. This likely involves either an on-chain calculation of days since a start date or an off-chain oracle pushing daily updates—either way, the piece is programmatically tied to the real-world passage of time and will continue to tick until Assange’s status changes. The open edition was dynamic in a different sense: when each Censored NFT was first minted, the text submitted by the user was stored (on-chain or via metadata) but visually represented in a redacted style (for example, as black bars or scrambled characters). Then, upon Assange’s release on 24 June 2024, a contract function executed to “uncensor” all these tokens, updating the metadata so that the true text became visible. Each open-edition NFT is thus unique in message but identical in having undergone the same transformation from obscured to revealed. The fundraising mechanism was straightforward: the sale smart contract directed the incoming Ether to designated wallets associated with Assange’s defense fund and other aligned charities (indeed, a portion of the 671 ETH raised from the open edition was later donated to support journalistic freedom and even humanitarian aid). The dynamics here combined artistic intent with real-world impact: the more people voiced themselves (minted NFTs) and the higher the DAO bid for Clock, the more tangible support was generated for the cause. In a sense, the blockchain functions as both canvas and conduit in Censored, ensuring that the art’s message (the imperative of free expression) is inseparable from the action it precipitates (material aid for a free press).
@@
The participatory element of Censored was striking. For Clock, participation took the form of collective fundraising: an Assange-supporter DAO (decentralized autonomous organization) pooled resources from over 10,000 people to bid on and ultimately win the piece for 16,593 ETH (about \$52.8 million), making each contributor a part-owner of the artwork and massively amplifying its political message. For the open edition, Pak opened a web portal on February 7, 2022, where anyone could type a short message and mint it as an NFT. There were minimal barriers—aside from blockchain transaction fees—which allowed thousands of ordinary people worldwide to inscribe their uncensored thoughts onto the blockchain. During the mint period, participants wrote messages ranging from political statements (“FREE ASSANGE NOW”) to personal reflections, fully aware that they would initially be recorded in a censored (illegible) state. After the mint closed, Pak triggered a global reveal: all the NFTs’ images were updated to show the actual text of the messages, symbolizing voices breaking through suppression. In this way, the audience’s role wasn’t just to observe but to speak and be memorialized as part of the art. Additionally, by buying these open-edition NFTs (each priced modestly), participants contributed to a fundraiser—all proceeds were directed to Assange’s legal defense and pro-freedom organizations—blending activism with artistic engagement. In essence, Censored turned its audience into an active assembly of publishers and protesters, demonstrating the principle that in a decentralized network, everyone can raise their voice and collectively bolster a cause.
@@
In Censored, Pak acted not only as an artist but also as a facilitator of collective expression and activism. Pak and Assange’s team set the stage—defining the parameters of the open edition (time frame, interface for input, price) and crafting the Clock NFT—but then handed the mike to the public. For the open edition, Pak exercised restraint in content moderation: applying no filters in alignment with the project’s absolutist free speech stance, Pak allowed minters to submit any text, which was a radical trust in the community. However, Pak retained control over the overall execution: Pak decided when the reveal would happen and ensured it occurred uniformly for everyone, and Pak controlled the smart contract that forwarded funds to the intended beneficiaries. With Clock, once it was launched, the control shifted to the AssangeDAO (as the new owners) in terms of display and eventual fate of the piece, although the daily increment feature was baked into the token itself. In essence, Pak’s control was about creating a secure and open channel for expression and then stepping back at the right moments. The success of Censored depended on this light-touch approach; too much control (e.g., censoring the messages or manipulating the outcomes) would have undermined the trust and authenticity the piece needed to resonate. By structuring it as Pak did, Pak underscored Pak’s role as an enabler—Pak controlled the framework but not the voices within, embodying the artwork’s ethos in the very way it was managed.
Collective Emergent Behavior
@@
The Censored project galvanized a uniquely blended community of art enthusiasts, cypherpunks, transparency advocates, and supporters of Assange. The formation of AssangeDAO itself was a landmark event: thousands of strangers coordinated in a matter of days on chat platforms, raised tens of millions in cryptocurrency, and collectively won the Clock auction, an unprecedented feat of decentralized crowd-funding. This feat gave participants a profound sense of accomplishment—each member of the DAO owns a fractional piece of Clock via tokens, making Clock a collectively owned symbol of protest. In parallel, the open edition saw 29,766 NFTs minted across its 48-hour run, reflecting broad participation from the crypto community and beyond. Upon the reveal, a new collective experience began: people scoured the now-uncensored messages, sharing poignant or powerful quotes they found among the NFTs. The collection became a digital chorus of statements about freedom, justice, and personal sentiments, and owning one meant being part of that chorus. A notable emergent behavior was the way traditional lines blurred: activists became art collectors by minting NFTs, and NFT collectors became activists by joining the political messaging. The Censored Discord and other forums buzzed with discussion not just about the art’s value, but about the cause it supported and the stories behind individual messages. This convergence of normally separate communities (art, crypto, activism) was itself a product of the project. In the aftermath, the collective energy didn’t dissipate immediately: the DAO refocused on continued advocacy for Assange, and many open-edition holders continued to promote the messages they had minted, effectively using their NFTs as badges of alignment with the free speech movement. Censored thus forged an alliance of convenience into a lasting community centered on principle and expression.
@@
Censored stands out as a poignant union of art and activism, transforming the cold mechanics of blockchain into an emotional and political narrative. The poetic substance of the work arises from its contrasts: a single, ever-ticking clock counting the days of one man’s silencing, and a sea of uncensorable messages giving voice to thousands. The Clock NFT is minimalistic yet profound—each passing day it displays is a testimony to injustice, a visual poem of waiting and resilience. The open edition, once uncensored, became an anthology of global voices; its poetry is both literal (many entries were written with genuine passion or wit) and metaphorical, in that it turned censorship on its head—what was once hidden is now immortalized in public view. There is a strong sentiment that Censored wasn’t just documenting a moment in crypto art, but a moment in history: it bridged the gap between the decentralized art world and real-world social issues, demonstrating that the blockchain community could rally around something deeply human. The broader impact of Censored lies in its demonstration of how digital art can be immediately socially relevant. It set a precedent for using NFTs as a medium of protest and fundraising, expanding the notion of what kind of statements can be made—and preserved—through art. In summary, Censored made poetic meaning through direct action: it is art as a verb, an event in which aesthetic expression, technological infrastructure, and moral conviction converged to powerful effect.
@@
The Clock NFT sold on 9 February 2022 for 16,593 ETH (approximately \$52.8 million), bought by AssangeDAO, a collective of 10,000+ members pooling funds. This made it one of the most expensive NFTs ever and provided a huge donation to Assange’s legal defense. The open edition X/X minted 29,766 NFTs over its 48-hour run, raising 671 ETH in voluntary contributions (about \$2.1 million) for pro-freedom organizations chosen by Pak and Assange. Each of these message NFTs was initially non-transferable; following Assange’s release on 24 June 2024, the tokens were unlocked and the censored messages revealed. Over 17,000 unique Ethereum addresses participated in the mint (many contributed multiple messages via additional wallets). The largest donation from a single minter was 50 ETH (showing some used the mint as a way to donate significantly). Post-drop, AssangeDAO has kept Clock on display via a fractionalized ownership governance token (representing membership shares in the DAO). Censored stands as a record-breaking instance of political fundraising through digital art, and a case where an NFT project directly engaged international press and communities far beyond the traditional art world.
Annotation
Censored is the most politically charged and arguably the most poetically powerful work in the portfolio, serving as the strongest case for C6 (poetic message embedding): the entire interaction ritual—submitting a message that is initially blacked out, then revealed only upon Assange’s release—enacts the experience of censorship and liberation in a way that purely visual art cannot. The political claim embedded in the work is inseparable from the collective behavior it demands: nearly 30,000 participants became co-authors of a collective “censored chorus,” each contributing a personal message that functioned simultaneously as artistic expression and political advocacy for Assange’s release (C3, distributed agency). The poetic message was literally embedded into the NFT, and the entire arc—from locked, non-transferable tokens bearing blacked-out text to their eventual unlocking and revelation upon Assange’s release on 24 June 2024—constitutes a profoundly poetic process with genuine geopolitical meaning. The non-transferability mechanism, inherited from the soul-bound logic pioneered in Hate / Invisible Mechanisms, here acquires a new resonance: the tokens were imprisoned alongside Assange, and their liberation mirrored his. C2 (autonomous governance) takes an unusual form: AssangeDAO’s collective ownership of Clock, and the reveal mechanism’s binding to an external real-world event, make the artwork’s governance partly extra-protocol—dependent on geopolitical outcomes beyond any participant’s control. This external dependency is unique in the portfolio and raises the question of whether protocol art’s autonomy can meaningfully extend to conditions outside the blockchain. The whole process—collective advocacy encoded in code, censorship enacted and then reversed, tokens locked and then freed—is deeply powerful, demonstrating that protocol art can achieve a form of political poetry in which the medium’s constraints and affordances carry moral weight.